# Pentagram maps over rings, Grassmannians, and skewers

### v2

@@ 3.5 The projective plane over the dual numbers
We note that, when restricted to lines through the origin, the map $\phi_{P}$ is essentially the identity map on $\mathbb{R}\mathbb{P}^{2}$, as it sends such a line to its direction; as for $\phi_{L}$, it is the polar duality map sending a line to its orthogonal plane. In this case, Proposition 3.21 just says that points $P_{1},P_{2}\in\mathbb{R}\mathbb{P}^{2}$ are polar if and only if $P_{1}$ belongs to the polar line of $P_{2}$.
@@ 4.1 Polygons in projective spaces over rings
Over a field, the pentagram map is well-defined on generic polygons. In particular, one can always apply the pentagram to a polygon such that all consecutive triples of vertices are in general position. The general position requirement means that the points are not collinear. In the context of (stably finite) rings, the right condition that ensures that the pentagram map is defined is the spanning condition.
@@
Let $M$ be an $R$-module. We say that points $P_{1},\dots,P_{n}\in\mathbb{P}(M)$ span $\mathbb{P}(M)$ if $P_{1}+\dots+P_{n}=M$.
@@
We notice that our definition of polygons allows either infinite polygons or closed ones, i.e., polygons such that $P_{i}=P_{i+n}$ for some fixed $n\in\mathbb{Z}_{+}$ and all $i\in\mathbb{Z}$.
@@
We say two polygons $(P_{i})$ and $(P_{i}^{\prime})$ are projectively equivalent if there exists a fixed projective transformation that takes each $P_{i}$ to $P_{i}^{\prime}$.
@@
We note that the notion of being spanning, and hence the notion of a polygon, is not determined by the incidence structure. For example, it is easy to see that the inclusion map $\mathbb{Z}^{3}\to\mathbb{Q}^{3}$ gives an isomorphism of incidence spaces $\mathbb{P}^{2}(\mathbb{Z})\simeq\mathbb{P}^{2}(\mathbb{Q})$ (the same is true for any PID and its field of fractions). However, while any three non-collinear points $P_{1},P_{2},P_{3}$ span $\mathbb{P}^{2}(\mathbb{Q})$, for $\mathbb{P}^{2}(\mathbb{Z})$ that is only so if the determinant of the matrix formed by the basis vectors in $P_{1},P_{2},P_{3}$ is ±$\pm 1$. Accordingly, not any polygon over $\mathbb{Q}$ is a polygon over $\mathbb{Z}$ in the sense of Definition 4.2, even though $\mathbb{P}^{2}(\mathbb{Z})\simeq\mathbb{P}^{2}(\mathbb{Q})$ as incidence spaces.
@@ 4.2 Pentagram maps over stably finite rings, Grassmannians, and skewers
First, observe that since $M=P_{2}\oplus(P_{3}\oplus P_{4})$, and $P_{2}\subset P_{1}\oplus P_{2}$, we have $P_{1}\oplus P_{2}=P_{2}\oplus((P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4}))$, and hence
$M=(P_{2}\oplus P_{3})\oplus((P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})).$
So, the submodule $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})$ admits a direct complement. Further, we also have
$M=(P_{2}\oplus P_{3})\oplus P_{1},$
so $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})\simeq P_{1}$, with the isomorphism given by the projection along $P_{2}\oplus P_{3}$. Thus, $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})$ is indeed a subspace of rank one, as needed. ∎
Applying Lemma 4.6 to vertices $P_{i},P_{i+2},P_{i+1},P_{i+3}$ (in that specific order) of a polygon $P_{i}$, we get that the pentagram map on polygons is well-defined. We note that the pentagram map is not a map from the set of polygons to itself, since the image of a polygon under the pentagram map is not necessarily a polygon. Instead, it should be understood as a binary relation on the set of polygons. In the case when the base ring is a field, the pentagram map is, actually, a rational mapping (for explicit formulas see, e.g., [15]).
@@
This is indeed the inverse of the pentagram map in the sense that the composition of this map and the pentagram map (taken in either order) is the identity, up to shifting indices, wherever that composition is defined.
We note that both the pentagram map and its inverse commute with projective transformation and, as such, are well-defined on the set of polygons modulo projective equivalence. Also, note that the image of a closed polygon under the pentagram map is again a closed polygon.
@@
The pentagram map over the matrix ring $\text{M}_{k}(\mathbb{F})$ is precisely (the planar case of) the Grassmann pentagram map, as defined in [4]:
@@
A polygon in $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ is a sequence of $k$-dimensional subspaces $V_{i}\subset\mathbb{F}^{3k}$ such that $V_{i}\oplus V_{i+1}\oplus V_{i+2}=\mathbb{F}^{3k}$. The pentagram map send such a sequence to the sequence
@@
We have $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))=\mathbb{P}(V^{k})$, where $V$ is a vector space over $\mathbb{F}$ of dimension $3k$. The correspondence $W\mapsto W^{k}$ identifies the lattice of subspaces of $V$ with lattice of submodules of $V^{k}$. Under this correspondence, points of $\mathbb{P}(V^{k})$ are identified with $k$-dimensional subspaces of $V$. Three points span $\mathbb{P}(V^{k})$ if and only if the corresponding $k$-dimensional subspaces span $V$. Thus, a polygon in $\mathbb{P}(V^{k})$ is precisely a collection $V_{i}\in\mathrm{Gr}(k,V)$ such that $V_{i}\oplus V_{i+1}\oplus V_{i+2}=V$. Further, intersections and spans of subspaces in $V^{k}$ correspond to intersections of subspaces in $V$, giving the claimed formula for the pentagram map. ∎
@@
The pentagram map over the ring $\mathbb{D}$ of dual numbers is the skewer pentagram map, as defined in [21]:
@@
Therefore, three affine lines span $\mathbb{P}^{2}(\mathbb{D})$ if and only if their direction vectors span $\mathbb{R}^{3}$, giving the desired description of polygons. As for the formula for the pentagram map, it follows from the definition of the latter and Proposition 3.21. ∎

### v3

@@ 3.5 The projective plane over the dual numbers
We note that, when restricted to lines through the origin, the map $\phi_{\mathrm{P}}$ is essentially the identity map on $\mathbb{R}\mathbb{P}^{2}$, as it sends such a line to its direction; as for $\phi_{\mathrm{L}}$, it is the polar duality map sending a line to its orthogonal plane. In this case, Proposition 3.23 just says that points $P_{1},P_{2}\in\mathbb{R}\mathbb{P}^{2}$ are polar if and only if $P_{1}$ belongs to the polar line of $P_{2}$.
Remark 3.25
[…]
3.6 Projective spaces over path algebras
In this section, we describe projective spaces over path algebras of quivers, with the algebra of triangular matrices as the main example. We will see that these can be understood in terms of flags and Schubert cells in the Grassmannian. For a detailed account of quivers and their path algebras, see, e.g., [13, 20].
Let $Q$ be a quiver, i.e., a directed graph, possibly with multiple edges and loops. Fix some ground field $\mathbb{F}$.
Definition 3.26
The path algebra $\mathbb{F}Q$ of $Q$ is an $\mathbb{F}$-algebra spanned, as an $\mathbb{F}$-vector space, by directed paths in $Q$, including, for every vertex $i$ of $Q$, a trivial path $e_{i}$, starting and ending at $i$. Multiplication in $\mathbb{F}Q$ corresponds to concatenation of paths: $ab$ is a path obtained by following first $b$ and then $a$. Concatenation $ab$ is only defined when the endpoint of $b$ coincides with the starting point of $a$. If concatenation is undefined, then the product of the corresponding paths is 0.
For any quiver $Q$, its path algebra $\mathbb{F}Q$ is unital and associative. The identity element is given by $1=\sum e_{i}$, where the sum is taken over all vertices $i$ of $Q$. The path algebra is finite-dimensional if and only if the quiver has no directed cycles.
Example 3.27
Consider a Dynkin quiver of type $A_{k}$, with all arrows pointing in one direction:
$1\to 2\to\dots\to k.$
For every $i\leq j$, where $1\leq i,j\leq k$, this quiver has a unique directed path from $i$ to $j$. Denote it by $E_{ji}$. Then the corresponding path algebra is spanned by $E_{ij}$, $i\geq j$, with multiplication given by $E_{ij}E_{jl}=E_{il}$ and all other products being 0. This is precisely the algebra $\mathrm{T}_{k}(\mathbb{F})$ of $k\times k$ lower-triangular matrices over $\mathbb{F}$, with $E_{ij}$ being matrix units, cf. [20, Example 4.3] (note that the algebras of upper- and lower-triangular matrices are isomorphic; we choose to work with lower-triangular matrices to simplify notation).
Modules over path algebras can be understood in terms of representations of the corresponding quiver.
Definition 3.28
A representation of a quiver $Q$ is an assignment of a vector space $V_{i}$ to its every vertex $i$ and a linear map $\phi_{a}\colon V_{i}\to V_{j}$ to every directed edge $a$ from $i$ to $j$.
Subrepresentations, isomorphisms, direct sums, etc. of quiver representations are defined in an obvious way.
There is a natural identification between quiver representations and (with our conventions, left) modules over the corresponding path algebra, see [13, Theorem 1.7]. Given a $\mathbb{F}Q$-module $M$, set $V_{i}:=e_{i}M$; these are the vector spaces assigned to the vertices of $Q$. The map $\phi_{a}\colon V_{i}\to V_{j}$ assigned to a directed arrow $a$ from $i$ to $j$ is given by the action of $a$, viewed as an element of $\mathbb{F}Q$. We have $aV_{i}\subset V_{j}$ since $ae_{i}=e_{j}a$.
Conversely, given a representation of a quiver $Q$, the corresponding $\mathbb{F}Q$-module is defined as the direct sum of vector spaces $V_{i}$ assigned to the vertices of $Q$. The action of $\mathbb{F}Q$ on $\bigoplus V_{i}$ is defined as follows: if $a$ is an arrow from vertex $i$ to vertex $j$, then the action of $a$ is given by $av=\phi_{a}(v)$ if $v\in V_{i}$, and $av=0$ otherwise.
Example 3.29
A representation of the quiver from Example 3.27 is a sequence of vector spaces and maps
$V_{1}$$V_{2}$$\dots$$V_{k-1}$$V_{k}$$\scriptstyle{\phi_{1}}$$\scriptstyle{\phi_{2}}$$\scriptstyle{\phi_{k-2}}$$\scriptstyle{\phi_{k-1}}$
The associated module over the algebra $\mathrm{T}_{k}(\mathbb{F})$ of lower-triangular matrices is the vector space $V_{1}\oplus\dots\oplus V_{k}$. The matrix $E_{ii}$ acts as a projector onto $V_{i}$. The matrix $E_{i+1,i}$ acts as $\phi_{i}$ on $V_{i}$ and trivially on other $V_{j}$.
The natural $\mathrm{T}_{k}(\mathbb{F})$-module structure on $\mathbb{F}^{k}$ corresponds to all $V_{i}$ being one-dimensional and all $\phi_{i}$ being isomorphisms, while $\mathrm{T}_{k}(\mathbb{F})$ as a module over itself corresponds to $\dim V_{i}=i$, and all $\phi_{i}$ being injective.
This interpretation of modules over path algebras makes the description of corresponding projective spaces a relatively straightforward task. From now on, we will concentrate on the quiver from Example 3.27 corresponding to the algebra $\mathrm{T}_{k}(\mathbb{F})$ of triangular matrices. We will see that the corresponding projective spaces can be understood in terms of flags. More general quivers lead to configurations of subspaces with more complicated combinatorics, see Example 3.33 below.
Consider a representation of the quiver from Example 3.27, as described in Example 3.29. Assuming that all maps $\phi_{i}$ are injective (which is the case for free modules), we can think of such a representation as a flag $V_{1}\subset\dots\subset V_{k}$. If we further assume that all $V_{i}$ are finite-dimensional (again, that is so for free modules), then such a flag is determined, up to isomorphism, by its type, i.e., a non-decreasing sequence of integers $\dim V_{1}\leq\dots\leq\dim V_{k}$. A free rank one module over $\mathrm{T}_{k}(\mathbb{F})$ corresponds to a complete flag, i.e., a flag of type $(1,\dots,k)$, see Example 3.29. So, given a flag, the projectivization of the corresponding $\mathrm{T}_{k}(\mathbb{F})$ module consists of its complete subflags which admit a direct complement. These are described by the following:
Proposition 3.30
Suppose $V_{1}\subset\dots\subset V_{k}$ is a flag and $U_{1}\subset\dots\subset U_{k}$ is its subflag, i.e., $U_{i}\subset V_{i}$. Then the following are equivalent:
  1. 1.
    The subflag $U_{1}\subset\dots\subset U_{k}$ admits a direct complement in $V_{1}\subset\dots\subset V_{k}$, i.e., there is another subflag $W_{1}\subset\dots\subset W_{k}$ such that $V_{i}=U_{i}\oplus W_{i}$ for all $i=1,\dots,k$.
  2. 2.
    $U_{i}=V_{i}\cap U_{i+1}$ for all $i=1,\dots,k-1$.
  3. 3.
    $U_{i}=V_{i}\cap U_{k}$ for all $i=1,\dots,k$.
Proof.
$1\implies 2$. We have $U_{i}\subset V_{i}\cap U_{i+1}$ since $U_{1}\subset\dots\subset U_{k}$ is a subflag of $V_{1}\subset\dots\subset V_{k}$. To show that $U_{i}\supset V_{i}\cap U_{i+1}$, suppose $v\in V_{i}\cap U_{i+1}$. Since $v\in V_{i}$ and $V_{i}=U_{i}\oplus W_{i}$, we have $v=u+w$ where $u\in U_{i}$, $w\in W_{i}$. Since $u\in U_{i}$, we also have $u\in U_{i+1}$. Since $v\in U_{i+1}$, it follows that $w\in U_{i+1}$. On other hand $w\in W_{i}\subset W_{i+1}$, and $U_{i+1}\cap W_{i+1}=0$, so $w=0$, and $v=u\in U_{i}$, as needed.
$2\implies 1$. We prove the result by induction on $k$. The statement is vacuously true for $k=1$. For general $k$, by induction hypothesis, the subflag $U_{1}\subset\dots\subset U_{k-1}$ of $V_{1}\subset\dots\subset V_{k-1}$ admits a complementary flag $W_{1}\subset\dots\subset W_{k-1}$. To extend it to a flag $W_{1}\subset\dots\subset W_{k}$ complementary to $U_{1}\subset\dots\subset U_{k}$, we need to find a subspace $W_{k}\supset W_{k-1}$ such that $V_{k}=U_{k}\oplus W_{k}$. Clearly, such a subspace exists as long as $U_{k}\cap W_{k-1}=0$. To see the latter, observe that $U_{k}\cap W_{k-1}\subset U_{k}\cap V_{k-1}=U_{k-1},$ which has a trivial intersection with $W_{k-1}$. Thus, $U_{k}\cap W_{k-1}=0$.
$2\implies 3$. Assume that $U_{i}=V_{i}\cap U_{i+1}$ for all $i=1,\dots,k-1$. Then, for any such $i$, we have
$U_{i}=V_{i}\cap U_{i+1}=V_{i}\cap(V_{i+1}\cap U_{i+2})=\dots=V_{i}\cap\dots\cap V_{k}\cap U_{k}=V_{i}\cap U_{k}.$
Also, $U_{i}=V_{i}\cap U_{k}$ is trivially true for $i=k$. The result follows.
$3\implies 2$. Assume that $U_{i}=V_{i}\cap U_{k}$. Then
$U_{i}=U_{i}\cap U_{i+1}=(V_{i}\cap U_{k})\cap U_{i+1}=V_{i}\cap(U_{k}\cap U_{i+1})=V_{i}\cap U_{i+1}.$
So, subflags of $V_{1}\subset\dots\subset V_{k}$ admitting a direct complement are those of the form $U\cap V_{i}$, where $U=U_{k}$ is a subspace of $V_{k}$. In particular, we obtain the following:
Proposition 3.31
Points in the projectivization of a $\mathrm{T}_{k}(\mathbb{F})$-module associated with a flag $V_{1}\subset\dots\subset V_{k}$ correspond to subspaces $U\in\mathrm{Gr}(k,V_{k})$ such that $\dim(U\cap V_{i})=i$ for all $i=1,\dots,k$, cf. Proposition 3.11.
It is easy to see that the set of such $U$ is a partial closure of a Schubert cell in $\mathrm{Gr}(k,V_{k})$ corresponding to a partition $\lambda_{i}:=(\dim V_{k}-k)-(\dim V_{i}-i)$, provided that the sequence $\lambda_{i}$ is non-increasing, i.e., $\dim V_{i+1}>\dim V_{i}$ for all $i$. If this condition does not hold, then a complete subflag of $V_{1}\subset\dots\subset V_{k}$ cannot admit a direct complement, so the projectivization of the corresponding module is empty.
Further, since free rank one $\mathrm{T}_{k}(\mathbb{F})$-modules correspond to complete flags, a free rank $n$ module corresponds to a flag of type $(n,2n,\dots,kn)$. In particular, we have the following:
Proposition 3.32
The projective plane $\mathbb{P}^{2}(\mathrm{T}_{k}(\mathbb{F}))$ consists of $P\in\mathrm{Gr}(k,3k)$ such that $\dim(P\cap\mathbb{F}^{3i})=i,$ where $\mathbb{F}^{3}\subset\dots\subset\mathbb{F}^{3n}$ is a flag of coordinates subspaces. Lines in that projective plane are given by $L\in\mathrm{Gr}(2k,3k)$ such that $\dim(L\cap\mathbb{F}^{i})=2i.$ Incidence between points and lines is given by containment, cf. Proposition 3.17.
For instance, points in $\mathbb{P}^{2}(\mathrm{T}_{2}(\mathbb{F}))$ are $2$-planes in $\mathbb{F}^{6}$ which meet a fixed $3$-plane along a line, while lines are $4$-planes which meet the same $3$-plane along a $2$-plane.
In a similar way, one can describe projective spaces for path algebras of more complicated quivers. In that case, instead of flags one needs to study more complex configurations of subspaces.
Example 3.33
Modules over the path algebra of the quiver $1\to 3\leftarrow 2$ correspond to triples of vector spaces $V_{1},V_{2},V_{3}$ with linear maps $\phi_{1}\colon V_{1}\to V_{3}$, $\phi_{2}\colon V_{2}\to V_{3}$. Assuming that $\phi_{1},\phi_{2}$ are injective and their images have trivial intersection (which is the case for free modules), this is the same as a vector space $V_{3}$ with two trivially intersecting subspaces $V_{1}$, $V_{2}$. For a free rank one module one has $\dim V_{1}=\dim V_{2}=1$, $\dim V_{3}=3$. Using this, it is easy to see that the projective plane over this algebra consists of $3$-planes $U\in\mathrm{Gr}(3,9)$ which meet each of the two given $3$-planes (which are transverse to each other) along a line.
@@ 4.1 Polygons in projective spaces over rings
Over a field, the pentagram map is well-defined for polygons such that all consecutive triples of vertices are in general linear position. The general linear position requirement means that the points are not collinear. In the context of (stably finite) rings, the right condition that ensures that the pentagram map is defined is the spanning condition.
@@
Let $M$ be an $R$-module. We say that points $P_{1},\dots,P_{n}\in\mathbb{P}(M)$ span $\mathbb{P}(M)$ if their corresponding subspaces span $M$: $P_{1}+\dots+P_{n}=M$.
@@
We notice that our definition of polygons allows for infinite polygons or closed ones, that is, polygons such that $P_{i}=P_{i+n}$ for some fixed $n\in\mathbb{Z}_{+}$ and all $i\in\mathbb{Z}$.
@@
We say two polygons $(P_{i})$ and $(P_{i}^{\prime})$ are projectively equivalent if there exists a projective transformation that takes each $P_{i}$ to $P_{i}^{\prime}$.
@@
We note that the notion of being spanning, and hence the notion of a polygon, is not determined by the incidence structure. For example, it is easy to see that the inclusion map $\mathbb{Z}^{3}\to\mathbb{Q}^{3}$ gives an isomorphism of incidence spaces $\mathbb{P}^{2}(\mathbb{Z})\simeq\mathbb{P}^{2}(\mathbb{Q})$ (the same is true for any PID and its field of fractions).
However, while any three non-collinear points $P_{1},P_{2},P_{3}$ span $\mathbb{P}^{2}(\mathbb{Q})$, for $\mathbb{P}^{2}(\mathbb{Z})$ that is only so if the determinant of the matrix formed by the basis vectors in $P_{1},P_{2},P_{3}$ is $\pm 1$. Accordingly, not every polygon over $\mathbb{Q}$ is a polygon over $\mathbb{Z}$ in the sense of Definition 4.2, even though $\mathbb{P}^{2}(\mathbb{Z})\simeq\mathbb{P}^{2}(\mathbb{Q})$ as incidence spaces.
@@ 4.2 Pentagram maps over stably finite rings, Grassmannians, and skewers
First, observe that, since $M=(P_{3}\oplus P_{4})\oplus P_{2}$, and $P_{2}\subset P_{1}\oplus P_{2}$, we have
$P_{1}\oplus P_{2}=((P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4}))\oplus P_{2}.$ (4.1)
So, there is an isomorphism of submodules $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})\simeq P_{1}$, given by projection along $P_{2}$. In particular, $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})$ is a rank one free module. Also, using (4.1), we get
$M=P_{1}\oplus P_{2}\oplus P_{3}=((P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4}))\oplus P_{2}\oplus P_{3}.$
So, the rank one free submodule $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})$ admits a direct complement in $M$, i.e., is a point. ∎
Applying Lemma 4.6 to vertices $P_{i},P_{i+2},P_{i+1},P_{i+3}$ (in that specific order) of a polygon $P_{i}$, we get that the pentagram map on polygons is well-defined. We note that the pentagram map is not a map from the set of polygons to itself, since the image of a polygon under the pentagram map is not necessarily a polygon. Instead, it should be understood as a binary relation on the set of polygons. In the case when the base ring is a field, the pentagram map is, actually, given by rational formulas (see, e.g., [18, Section 2.2]).
@@
This is indeed the inverse of the pentagram map in the sense that the composition of this map and the pentagram map (taken in either order) is the identity, up to shifting indices, wherever that composition is defined. To see this, consider a polygon $P_{i}\in\mathbb{P}^{2}(R)$ that is mapped by the pentagram map to a polygon $P^{\prime}_{i}\in\mathbb{P}^{2}(R)$. Then, by construction, we have that $P_{i}^{\prime},P_{i+1}^{\prime}$ belongs to the line $P_{i+1}P_{i+3}$. Since $P_{i}^{\prime}$ is a polygon, it follows that $P_{i}^{\prime}P_{i+1}^{\prime}=P_{i+1}P_{i+3}$ and $P_{i}^{\prime}P_{i+1}^{\prime}\cap P_{i+2}^{\prime}P_{i+3}^{\prime}=P_{i+3}$. Thus, the pentagram map followed by the inverse pentagram is the identity (up to shifting indices). Likewise, if we apply the inverse map first, we get the identity as well.
We note that both the pentagram map and its inverse commute with projective transformations and, as such, are well-defined on the set of polygons modulo projective equivalence. Also, note that if the image of a closed polygon under the pentagram map is again a polygon, then it is also closed.
@@
The pentagram map over the matrix ring $\mathrm{M}_{k}(\mathbb{F})$ is precisely (the planar case of) the Grassmann pentagram map, as defined in [5, Definition 4.1]:
@@
A polygon in $\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F}))$ is a sequence of $k$-dimensional subspaces $V_{i}\subset\mathbb{F}^{3k}$ such that $V_{i}\oplus V_{i+1}\oplus V_{i+2}=\mathbb{F}^{3k}$. The pentagram map sends such a sequence to the sequence
@@
We have $\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F}))=\mathbb{P}(V^{k})$, where $V$ is a vector space over $\mathbb{F}$ of dimension $3k$. The correspondence $W\mapsto W^{k}$ identifies the lattice of subspaces of $V$ with the lattice of submodules of $V^{k}$. Under this correspondence, points of $\mathbb{P}(V^{k})$ are identified with $k$-dimensional subspaces of $V$. Three points span $\mathbb{P}(V^{k})$ if and only if the corresponding $k$-dimensional subspaces span $V$. Thus, a polygon in $\mathbb{P}(V^{k})$ is precisely a collection $V_{i}\in\mathrm{Gr}(k,V)$ such that $V_{i}\oplus V_{i+1}\oplus V_{i+2}=V$. Further, intersections and spans of subspaces in $V^{k}$ correspond to intersections of subspaces in $V$, giving the claimed formula for the pentagram map. ∎
Polygons and the pentagram map for triangular matrices have a similar description. Polygons in $\mathbb{P}^{2}(\mathrm{T}_{k}(\mathbb{F}))$ consist of sequences of subspaces $V_{i}\subset\mathbb{F}^{3k}$ such that, for all $i\in\mathbb{Z}$ and all $j=1,\dots,k$, we have
$\dim(V_{i}\cap\mathbb{F}^{3j})=j,\quad(V_{i}\cap\mathbb{F}^{3j})\oplus(V_{i+1}\cap\mathbb{F}^{3j})\oplus(V_{i+2}\cap\mathbb{F}^{3j})=\mathbb{F}^{3j}.$
The pentagram map on these sequences is given by the same formula as for the full matrix ring.
For example, a polygon in $\mathbb{P}^{2}(\mathrm{T}_{2}(\mathbb{F}))$ is a sequence of $2$-planes $V_{i}$ in $\mathbb{F}^{6}$ such that $V_{i}\oplus V_{i+1}\oplus V_{i+2}=\mathbb{F}^{6}$, and each $V_{i}$ intersects a given subspace $\mathbb{F}^{3}\subset\mathbb{F}^{6}$ along a line $W_{i}$. Additionally, $W_{i}\oplus W_{i+1}\oplus W_{i+2}=\mathbb{F}^{3}$. The pentagram map produces new sequences $V_{i}^{\prime},W_{i}^{\prime}$ given by
$V_{i}^{\prime}:=(V_{i}\oplus V_{i+2})\cap(V_{i+1}\oplus V_{i+3}),$ (4.3)
$W_{i}^{\prime}:=(W_{i}\oplus W_{i+2})\cap(W_{i+1}\oplus W_{i+3}).$ (4.4)
Notice that $W_{i}$, which are just points in $\mathbb{P}^{2}(\mathbb{F})$, undergo the usual pentagram map. Thus, the pentagram map in $\mathbb{P}^{2}(\mathrm{T}_{2}(\mathbb{F}))$ can be thought of as a Grassmannian pentagram map on sequences $V_{i}\in\mathrm{Gr}(2,6)$ coupled with the usual pentagram map on sequences $W_{i}\in\mathrm{Gr}(1,3)$. The coupling condition is $W_{i}=V_{i}\cap\mathbb{F}^{3}$. A similar description can be given for pentagram maps over $\mathrm{T}_{k}(\mathbb{F})$ or, more generally, arbitrary path algebras corresponding to acyclic quivers.
@@
The pentagram map over the ring $\mathbb{D}$ of dual numbers is the skewer pentagram map, as defined in [25, Section 9.7]:
@@
Therefore, three affine lines span $\mathbb{P}^{2}(\mathbb{D})$ if and only if their direction vectors span $\mathbb{R}^{3}$, giving the desired description of polygons.
As for the formula for the pentagram map, it follows from the definition of the latter and Proposition 3.23. ∎